OpenStudy (anonymous):

FUN!!! (not) Word problem: A student works two part time jobs. Job A he gets paid $8.00 per hour and Job B he gets pd $15.00 per hour. He ONLY works 25hours per week. This weeks he brings home $270.00. How long did he work each job?

OpenStudy (anonymous):

So call x the number of hours of job a and y the number of hours of job b "He ONLY works 25hours per week" x+y=25 "Job A he gets paid $8.00 per hour and Job B he gets pd $15.00 per hour, This week he brings home $270.00." 8x+15y=270 so any idea how to solve the simultaneous equations? x+y=25 8x+15y=270

OpenStudy (anonymous):

No... I didn't even realize it would be a double equation till you pointed it out (BTW.. good call and thank you cause I probably would have ended this session in tears.. lol)

OpenStudy (anonymous):

Hmm, well, do you know what substitution method is?

OpenStudy (anonymous):

I don't... I had a 'problem earlier, trying to grasp this concept but the gal confused me to the point of tears (hence my last statement about tears) ~ right now, my brain is so jumbled that nothing makes much sense.

OpenStudy (anonymous):

Well ill try to explain. substitution method is where you solve for a variable in terms of another and then use that expression to plug into the other equation an example: x+2y=4 3x+y=9 you can solve for x in the first equation x=-2y+4 now since x IS -2y+4 you can take that, and plug it into the second equation where x is 3(-2y+4)+y=9 now theres only one variable to solve for =D so now in our problem, apply the same concept, can you solve for x in the first equation?

OpenStudy (anonymous):

Is the answer... 16hr

OpenStudy (anonymous):

16@ $8.00 and 9 @ $15.00

OpenStudy (anonymous):

I just divided 270 by 2 = $135.0 0 per job. I divided $135. by 8 and $135 by 15

OpenStudy (anonymous):

Noo =/ try solving for x in the first equation